OpenStudy (anonymous):

Find a polynominal of degree 4with -2 as a zero of multiplicity 3 and 0 as a zero of multiplicity 1

OpenStudy (perl):

y = ( x - (- 2)) ^3 ( x - 0) y = ( x + 2)^3 * x

OpenStudy (perl):

y = (x + 2)(x+2)(x+2) x then expand that

OpenStudy (anonymous):

@perl can you come back to my question?

OpenStudy (perl):

ok

OpenStudy (perl):

tag me in question